# Generative AI and Firm Productivity:

### v1

@@ Appendix D Details on the Heterogeneous Treatment Effects Across Seller
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
Table D2 exhibited the heterogeneous impact of Marekting Push Message. Similarly, we observe that small sellers experience a significant increase in conversion rates from treated consumers (Columns 2, 4 and 6 in Panel B), while the conversion rates on big sellers show no significant changes between treatment and control groups. Thus, GenAI helps the platform to more effectively leverage content personalization in its marketing activities, particularly benefiting small sellers.
@@ Appendix E Details on the Heterogeneous Treatment Effects Across Consumers
In this experiment, we focus on the comparison between consumers who are served only by GenAI chatbot with those who are served by an auto-response indicating service unavailability. Table E1 reports the heterogeneous treatment effects across consumer groups. Our analysis reveals that the increases of both sales (Panel A) and converstion rate (Panel B) are more pronounced among newer consumers with shorter registration histories (Column 2), less active consumers with fewer logins days (Column 4), and lighter consumers with less past spending (Column 6).
@@
    We classify consumers into high and low groups and capture their online shopping experience using three metrics. Consumers in the low group are considered relatively inexperienced, as determined by meeting the following pre-experiment criteria: (1) a registration duration below the median among all consumer on the platform; (2) a number of 30-day login days below the platform-wide median; or (3) a 30-day total purchase amount below the platform-wide median.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
In Table E3, we observe similar pattern. There is significant and more pronounced sales increases for inexperienced consumers (Columns 2, 4 and 6). Augmenting product descriptions with GenAI-generated content substantially enhanced the sufficiency and clarity of product information, thereby motivating consumer purchase decisions, particularly among less experienced consumers.
@@ Appendix F Details on the Heterogeneous Treatment Effects Across Products
    We classify products into high and low groups based on three key pre-experiment dimensions. (1) Category market concentration, measured by the sales share of the top 1% of products (ranked by annual sales) within each category. Products in the low group belong to categories with concentration levels below the platform average. (2) Annual sales quantity, also defined within category. Products in the low group—referred to as tail products—are those comprising the bottom 50% cumulative share of total units sold when ranking products by annual sales quantity within each category. (3) Product price, defined within each category to control for category-level pricing variation. Low-priced products are those priced below the median of their respective category.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experiment period, and 0 otherwise.
@@
In Table F3, for market concentration and price level, we only observe significant gains for high-concentration categories (Column 1) and high-priced products (Column 5). The conversion rate increase is only significant for head products while the sales increase is significant for both head and tail products (Columns 3-4).

### v2

@@ Appendix D Details on the Heterogeneous Treatment Effects Across Sellers
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
Table D2 exhibited the heterogeneous impact of Marketing Push Message. Similarly, we observe that small sellers experience a significant increase in conversion rates from treated consumers (Columns 2, 4 and 6 in Panel B), while the conversion rates on big sellers show no significant changes between treatment and control groups. Thus, GenAI helps the platform to more effectively leverage content personalization in its marketing activities, particularly benefiting small sellers.
@@ Appendix E Details on the Heterogeneous Treatment Effects Across Consumers
In this experiment, we focus on the comparison between consumers who are served only by the GenAI chatbot and those who are served by an auto-response indicating service unavailability. Table E1 reports the heterogeneous treatment effects across consumer groups. Our analysis reveals that the increases of both sales (Panel A) and conversion rate (Panel B) are more pronounced among newer consumers with shorter registration histories (Column 2), less active consumers with fewer login days (Column 4), and lighter consumers with less past spending (Column 6).
@@
    We classify consumers into high and low groups and capture their online shopping experience using three metrics. Consumers in the low group are considered relatively inexperienced, as determined by meeting the following pre-experiment criteria: (1) a registration duration below the platform-wide median; (2) a number of 30-day login days below the platform-wide median; or (3) a 30-day total purchase amount below the platform-wide median.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
In Table E3, we observe similar pattern. There is a significant and more pronounced sales increase for inexperienced consumers (Columns 2, 4 and 6). Augmenting product descriptions with GenAI-generated content substantially enhanced the sufficiency and clarity of product information, thereby motivating consumer purchase decisions, particularly among less experienced consumers.
@@ Appendix F Details on the Heterogeneous Treatment Effects Across Products
    We classify products into high and low groups based on three key pre-experiment dimensions. (1) Category market concentration, measured by the sales share of the top 1% of products (ranked by annual sales) within each category. Products in the low group belong to categories with concentration levels below the platform average. (2) Annual sales quantity, also defined within the category. Products in the low group—referred to as tail products—are those comprising the bottom 50% cumulative share of total units sold when ranking products by annual sales quantity within each category. (3) Product price, defined within each category to control for category-level pricing variation. Low-priced products are those priced below the median of their respective category.
@@
    “Sales” represents the total expenditure on product orders. “Conversion Rate” measures consumers’ likelihood of making purchases. It is a binary indicator for purchase, which equals 1 if a consumer makes at least one order during the experimental period, and 0 otherwise.
@@
In Table F3, for market concentration and price level, we only observe significant gains for high-concentration categories (Column 1) and high-priced products (Column 5). The conversion rate increase is only significant for head products, while the sales increase is significant for both head and tail products (Columns 3-4).